who is faster? mike driving at 20.0m/s or john driving at 20.0km/h


1
Expert's answer
2021-09-07T14:07:21-0400

"v_{mike}=20\frac{m}{s}"

"v_{john}= 20\frac{km}{h}"

"1km = 1000m; 1h=3600s"

"v_{john}= 20\frac{km}{h}=\frac{20*1000}{3600}\frac{m}{s}\approx5.56\frac{m}{s}"

"v_{john}<v_{mike}"


"\text{Answer: Mike moves faster than John. }"